• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 23:59:16 on July28,2003

    Entry number 107794

    keyword=Target Shift Summary

    Target Log   Swing shift   July 28 2003

    D.S. Armstrong

    Shift Start:

    holding field: 0°
    target HWP: out
    lasers: 4,6,7 on
    Target: Pol 3He
    Cell: Priapus
    Beam: waiting for beam

       At start of shift, waiting for beam for harp/OTR/spot etc.
    checks after beam energy change... empty target position.

       Beam now looks good.

    18:25   Ask for beam off, go to reference
    cell position. Ask for 10 µA.

    18:38   Finished 10 min. run; now fill reference cell
    with 140 psig N2 and start another 10 minute run,
    this time at 2 µA.

    18:55   Finished nitrogen run; ask for target motion
    to go to empty cell so we can do a Moller measurement.

       Meanwhile, vent and evacuate Reference cell.

    21:40   Moller measurement completed. Dp spot++ check.
    21:53   Move to 3He cell
    21:58   do NMR measurement in 3He position.
    21:58    NMR measurement done, lasers on, but
    we are in the 3He cell position, not the pickup coils
    position (see HAlog entry


           Result: NMR polarization = 39.28%

        Leave target as/is, ask for beam back.

    23:30   Turn off beam to rotate to 270°. Turn off
    Lasers 4,6,7. Start rotation 0° to 270°.

    23:34   Turn on Laser 1

    23:35   Turn on Laser 2

    23:36   Turn on Laser 3

    Shift End:

    holding field: 270 °
    target HWP: out
    lasers: transverse: 1,2,3 ON
    Target: Pol 3He
    Cell: Priapus
    Beam: 3.0 µA

    hand off to owl shift, go home to sleep...